# Learning Patient-Specific Disease Dynamics with Latent Flow Matching for Longitudinal Imaging Generation

### v1

@@ Appendix A Training Setting
We adopt the AdamW optimizer (Loshchilov & Hutter, 2017) for all experiments. For the autoencoder (AE) and 3D U-Net (Ronneberger et al., 2015), we trained with a learning rate of $1\times 10^{-3}$ and a batch size of $2$ for 300 epochs. For the flow matching model based on the 3D U-Net, we used a reduced learning rate of $3\times 10^{-5}$ with a larger batch size of $4$ for 200 epochs.
@@ Appendix B Conditional Generation
Following Litrico et al. (Litrico et al., 2024), we encode the current sample time in flow matching into continuous embeddings using sinusoidal positional encodings. These time embeddings act as an additive bias to the input, ensuring that the generative process is aware of its temporal position.
Building on this, and inspired by BrLP (Puglisi et al., 2024), we introduce explicit control signals that combine patient-specific attributes with temporal queries. In particular, we condition the model not only on the start time, the current sample time, and the end query time, but also on the patient attributes. This design enables the model to incorporate both subject-specific and task-specific information when shaping the generative trajectory.
To inject these conditional signals into the U-Net backbone, we explored three strategies: 1. Additive biasing, 2. Cross-attention (Vaswani et al., 2017) and 3. Adaptive Layer Normalization (AdaLN) (Raffel et al., 2020).
@@ Appendix C Auto-encoder Ablation
We do not use deeper or wider AEs, as these configurations lead to out-of-memory (OOM) errors on our RTX A6000 GPU (48 GB). Based on the above results, we use channels $[64,128,256]$ and crop size $64^{3}$ as our default configuration. Larger crops (e.g., $72^{3}$) also result in OOM on our hardware. To keep the comparison with baselines fair, we do not use a pretrained AE, although initializing with MAISI Guo et al. (2025) can further improve PSNR (from 30.04 to 31.97 for $[64,128,256]$).
@@ Appendix I Discussion of LLM usage
We used large language models (LLMs), including ChatGPT (OpenAI, 2023) and Gemini (Team et al., 2023), as writing assistants for grammar correction and style refinement. All technical content, methodology, and experiments were developed by the authors.

### v2

@@ Appendix A Training Setting
We adopt the AdamW optimizer (loshchilov2017decoupled) for all experiments. For the autoencoder (AE) and 3D U-Net (ronneberger2015u), we trained with a learning rate of $1\times 10^{-3}$ and a batch size of $2$ for 300 epochs. For the flow matching model based on the 3D U-Net, we used a reduced learning rate of $3\times 10^{-5}$ with a larger batch size of $4$ for 200 epochs.
@@ Appendix B Conditional Generation
Following Litrico et al. (litrico2024tadm), we encode the current sample time in flow matching into continuous embeddings using sinusoidal positional encodings. These time embeddings act as an additive bias to the input, ensuring that the generative process is aware of its temporal position.
Building on this, and inspired by BrLP (puglisi2024enhancing), we introduce explicit control signals that combine patient-specific attributes with temporal queries. In particular, we condition the model not only on the start time, the current sample time, and the end query time, but also on the patient attributes. This design enables the model to incorporate both subject-specific and task-specific information when shaping the generative trajectory.
To inject these conditional signals into the U-Net backbone, we explored three strategies: 1. Additive biasing, 2. Cross-attention (vaswani2017attention) and 3. Adaptive Layer Normalization (AdaLN) (raffel2020exploring).
@@ Appendix C Auto-encoder Ablation
We do not use deeper or wider AEs, as these configurations lead to out-of-memory (OOM) errors on our RTX A6000 GPU (48 GB). Based on the above results, we use channels $[64,128,256]$ and crop size $64^{3}$ as our default configuration. Larger crops (e.g., $72^{3}$) also result in OOM on our hardware. To keep the comparison with baselines fair, we do not use a pretrained AE, although initializing with MAISI guo2025maisi can further improve PSNR (from 30.04 to 31.97 for $[64,128,256]$).
@@ Appendix I Discussion of LLM usage
We used large language models (LLMs), including ChatGPT (openai2023chatgpt) and Gemini (team2023gemini), as writing assistants for grammar correction and style refinement. All technical content, methodology, and experiments were developed by the authors.